Jim has 5 pieces of string. He needs to choose the piece that will be able to go around his 36-inch waist. His belt broke, and his pants are falling down. The piece needs to be at least 4 inches longer than his waist so he can tie a knot in it, but it cannot be more that 6 inches longer so that the ends will not show from under his shirt. Which of the following pieces of string will work the best?

3 4/5 feet
3 2/3 feet
3 3/8 feet
3 ¼ feet
2 ½ feet

Answer by rapaljer(4671)

Oh, I see what you mean! Jim is NOT using all 5 pieces of string--just the one that fits him the best.

The 3 feet exactly equals what Jim needs to go around his waist. He must have at least 4 inches (or 4/12 = 1/3 foot) more, but not more than 10 inches (or 10/12 = 5/6 feet) more.

Of your choices, only the "3 2/3" feet length fits between "3 1/3" and "3 5/6." The others are either too long or too short.

R^2 at SCC